the difference between two angles of a right angles triangle is 2 pi/5 radians . express angles in radians
 

Grade:11

2 Answers

Advait Nitin Amrutkar
8 years ago
the two angles are pi/ 20 and 9pi/ 20 radians.
 
let one angle be x.
so the other angle is x + 2 pi/ 5.
 
we know, x + ( x+ 2pi/ 5 ) = pi /2
solving we get x = pi/ 20 and  x + 2pi/5 = 9pi/20
grenade
8 years ago
pi/ 20 and 9pi/ 20 radian